### v1

@@ 5.2. Fibered domain of discontinuity
We first apply Theorem 5.4 which shows that $\rho$ is $\{n\}$-Anosov. Let $x\in\widetilde{S_{g}}$, we prove that the homotopy type of the boundary curve $\xi^{n}_{\rho}$ is the same as the homotopy type $[\xi_{u(x)}]$ from Proposition 6.3. We consider the map $\zeta_{\infty}$ that associates to $[q]\in\mathbb{S}u^{*}\mathcal{E}_{x}$ the limit of $\Phi_{t}([q])$ in $\partial\Gamma$. This map is homotopic in $\widetilde{S_{g}}\cup\partial\Gamma\setminus\{x\}$ to the map $\zeta_{1}$ that associates the projection $p(\Phi_{1})$. Hence by Lemma 4.3, $\zeta_{1}$ defines a diffeomorphism of degree $1$ between the circle $\mathbb{S}u(x)$ with its maximal orientation and the boundary $\partial\Gamma$ for the induced orientation.
The map $\xi^{n}_{\rho}\circ\zeta_{\infty}$ has the homotopy type $[\xi_{u(x)}]$ associated to $u(x)$: it defines a boundary map as in Proposition 6.3. The fact that $\rho$ is maximal is then a consequence of the characterization of maximal representations from Theorem 2.5. Indeed $u(x)\in\operatorname{Gr}_{2}^{\max}(\mathcal{Q})$ implies that $[\xi_{u(x)}]=n[\tau]$, and we already know that the degree of $\zeta_{\infty}$ is equal to the degree of $\zeta_{1}$ which is equal to $1$.
@@ 6.3. Construction of a fitting flow
Let $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$ be a maximal quadruple of Lagrangians, i.e. such that each cyclic oriented subtriple is maximal. The bilinear form $q_{\ell_{4},\ell_{3}}-q_{\ell_{1},\ell_{2}}$ is positive. In particular the zero set of these quadrics define two disjoint quadric hypersurfaces in $\mathbb{P}(\mathbb{R}^{2n})$.
@@
Let ${v}\in\mathbb{R}^{2n}$, it can be decomposed uniquely as ${v}={v}_{1}+{v}_{2}$ with ${v}_{1}\in\ell_{1}$ and ${v}_{2}\in\ell_{2}$. Moreover there exist some unique $x,y\in\ell_{4}$ such that ${v}_{1}=x+u_{1}(x)$ and ${v}_{2}=y+u_{2}(y)$. The vector ${v}$ decomposes therefore as ${v}=x+y+u_{1}(x)+u_{2}(y)$. One computes that :
@@
Finally the fact that $(\ell_{1},\ell_{2},\ell_{3})$ forms a maximal triple implies that the bilinear form $\omega\left(\cdot,(u_{2}-u_{1})(\cdot)\right)$ is positive. Therefore the previous expression, for $y\neq 0$, is strictly greater than $\omega\left(x,u_{1}(x)\right)+\omega\left(y,u_{1}(y)\right)+2\omega\left(y,u_{1}(x)\right)$, which is non-negative since $\omega\left(\cdot,u_{1}(\cdot)\right)$ is positive. In the case when $y=0$, this last inequality is strict for $x\neq 0$. Otherwise the previous inequality $\omega\left(y,(u_{2}-u_{1})(y)\right)>0$ is strict. Therefore for ${v}\neq 0$, $q_{\ell_{4},\ell_{3}}({v},{v})-q_{\ell_{1},\ell_{2}}({v},{v})>0$. ∎
@@
Let $\ell^{+},\ell^{-}:[0,1]\to\mathcal{L}_{n}$ be smooth and such that $\ell_{0}^{+}=\ell^{+}(0)$ and $\ell_{0}^{-}=\ell^{-}(0)$ are transverse and the linear maps $\dot{u}^{+}\in\operatorname{Hom}(\ell^{+}_{0},\ell^{-}_{0})$ and $\dot{u}^{-}\in\operatorname{Hom}(\ell^{+}_{0},\ell^{+}_{0})$ corresponding to $(\ell^{+})^{\prime}(0)$ and $(\ell^{-})^{\prime}(0)$ are such that $\omega(\cdot,\dot{u}^{+}(\cdot))$ and $\omega(\dot{u}^{-}(\cdot),\cdot)$ are positive respectively on $\ell^{+}_{0}$ and $\ell^{-}_{0}$. The derivative at $t=0$ of $q_{\ell^{+}(t),\ell^{-}(t)}$ is positive.
@@
The proof is similar to the previous one. Let $u^{+}(t)\in\operatorname{Hom}(\ell^{+}_{0},\ell^{-}_{0})$ and $u^{-}(t)\in\operatorname{Hom}(\ell^{-}_{0},\ell^{+}_{0})$ be the linear maps whose graph is equal to $\ell^{+}(t)$ and $\ell^{-}(t)$ respectively. Let ${v}\in\mathbb{R}^{2n}$. The derivative of the evaluation of $q_{t}=q_{\ell^{+}(t),\ell^{-}(t)}$ to ${v}$ can be written in term of the derivative $\dot{{v}}_{t}^{+}$, $\dot{{v}}_{t}^{-}$ of the vectors ${v}_{t}^{+}\in\ell_{+}(t)$, ${v}_{t}^{-}\in\ell_{-}(t)$ such that ${v}={v}_{t}^{+}+{v}_{t}^{-}$:
@@
$v^{\prime}$$\phi(w)$$y$$\zeta_{-J\phi(w)}$$\zeta_{J\phi(w)}$$x$$v$$w$$\zeta_{v}$$\zeta_{-Jw}$$\zeta_{Jw}$
@@
Let $w\in T^{1}_{x}\widetilde{S}$ be such that $\langle w,v\rangle>0$. The following quadruple is positive $(\zeta_{-Jw},\zeta_{-J\phi(w)},\zeta_{J\phi(w)},\zeta_{Jw})$ due to the negative curvature of the metric we put on $\widetilde{S_{g}}$, see Figure 4. Hence the corresponding quadruple of Lagrangians is maximal. Therefore $q_{\phi(w)}^{\circ}-q_{w}^{\circ}$ is positive by Lemma 6.11. When $\langle w,v\rangle<0$, the following quadruple is positive $(\zeta_{-J\phi(w)},\zeta_{-Jw},\zeta_{Jw},\zeta_{J\phi(w)})$ and therefore $q_{\phi(w)}^{\circ}-q_{w}^{\circ}$ is negative. Hence :
@@
Given an immersion $h:M\to\mathbb{P}(S^{2}V^{>0})$ from a manifold $M$ of dimension $d$ we define its Gauss map $\mathcal{G}h:M\to\operatorname{Gr}_{d}(\mathcal{Q})$, that associates to $x\in M$ the pencil $P$ associated with the codimension $d$ projective subspace of $\text{Gr}_{d}(\mathcal{Q})$ orthogonal to $u(h)$ at $h(x)$, for the $\operatorname{SL}(V)$-invariant Riemannian metric on $\mathbb{P}(S^{2}V)$.
@@
We first identify $\mathbb{P}(S^{2}V^{>0})$ with $\mathbb{P}(S^{2}(V^{*})^{>0})$ via the map $[X]\mapsto[X^{-1}]$. Note as once again we view elements of $S^{2}V$ and $S^{2}(V^{*})=\mathcal{Q}$ respectively as maps $V^{*}\to V$ and $V\to V^{*}$. An immersion $h:M\to\mathbb{P}(S^{2}V^{>0})$ hence defines a dual immersion $h^{*}:M\to\mathbb{P}(S^{2}(V^{*})^{>0})$. Fixing a volume form on $V$ and $V^{*}$ allows us to lift this map to a map $\overline{h}^{*}$ into the space of elements in $S^{2}(V^{*})^{>0}$ whose corresponding map $V\to V^{*}$ has determinant $1$. We define $\mathcal{G}h(x)=\mathrm{d}\overline{h}^{*}(T_{x}M)\subset S^{2}V^{*}=\mathcal{Q}$.
@@
Let $\gamma:\mathbb{R}\to M$ be a geodesic for the metric induced by $h$. We write the representative of $h(\gamma(o))$ with determinant $1$ in a basis $(e_{i})_{i\in I}$ such that for some $\lambda^{i}\in\mathbb{R}$, for all $t\in\mathbb{R}$:
@@
Let us fix a symplectic form $\omega$ on $V=\mathbb{R}^{2n}$. Let $\mathbb{X}_{\operatorname{Sp}}$ be the subset of $\mathbb{P}(S^{2}V^{>0})$ consisting of tensors $[q^{-1}]$ that are compatible with $\omega$, i.e. such that for some complex structure $J$ on $\mathbb{R}^{2n}$, $q+i\omega$ is a hermitian metric on $V$. Recall that $q:V\to V^{*}$ is a bilinear form, and $q^{-1}:V^{*}\to V$ is a tensor.
The space $\mathbb{X}_{\operatorname{Sp}}$ is a copy of the symmetric space associated to $\operatorname{Sp}(2n,\mathbb{R})$, which is a totally geodesic subspace of the symmetric space associated to the Lie group $\operatorname{SL}(2n,\mathbb{R})$ whose model is $\mathbb{P}(S^{2}V^{>0})$. However it is not a projective subset : the closure of the projective convex hull of $\mathbb{X}_{\operatorname{Sp}}$ in $\mathbb{P}(S^{2}V^{>0})$ is equal to $\mathbb{P}(S^{2}V^{\geq 0})$ since it contains all the extremal points of $\mathbb{P}(S^{2}V^{\geq 0})$, i.e. the rank one elements $S^{2}\mathbb{P}(\mathbb{R}^{2n})$.
@@
The annihilator of the tangent space to $\mathbb{X}_{\operatorname{Sp}}$ at this point can be identified with the space of following symmetric matrices where $A$ is symmetric and $B$ is antisymmetric:
@@
Suppose that the intersection is not transverse, i.e. that $q$ can be written in this form. Let $\ell^{+}$ and $\ell^{-}$ be Lagrangians on which $q$ is respectively positive and negative. Since the maximal compact $U(n)$ acts transitively on the space of Lagrangians one can assume that $\ell^{+}=\langle x_{1},\cdots,x_{n}\rangle$. Let $\ell^{-}=\langle x_{1},\cdots,x_{n}\rangle$. The Lagrangian $\ell^{-}$ is transverse to $\ell^{+}$ so for some symmetric matrix $U$, one can write this Lagrangian as the image of :
@@
This bilinear form is regular if and only if all the $\lambda_{i}$ are non-zero. If this is the case that it is positive on the Lagrangians $\langle x_{1},\cdots,x_{n}\rangle$ and negative on the Lagrangian $\langle y_{1},\cdots,y_{n}\rangle$. ∎
@@ Appendix A Immersions in the space of geodesics in $\mathbb{H}^{3}$
In this section we fix $V=\mathbb{C}^{2}$ and restrict ourselves to pencils of Hermitian quadrics. The result of this Section serve as an illustration of the previously introduced notions of fitting pairs and fitting directions.
@@
Let $S^{2}V=S^{2}_{h}V\oplus S^{2}_{a}V$ be the eigenspace decomposition for the operator $J\otimes J$ where $J$ is the complex conjugation. Here $S^{2}_{h}V$ is the $4$-dimensional eigenspace associated to $1$ and $S^{2}_{a}V=\mathcal{H}^{\circ}$, the $6$-dimensional space associated to $-1$. The intersection of $\mathbb{P}(S^{2}_{h}V)$ with the space of positive tensors is the projective Klein model for $\mathbb{H}^{3}$.
The annihilator of a pencil $P\in\operatorname{Gr}_{2}^{{\operatorname{mix}}}(\mathcal{H})$ is equal to $P^{\circ}=\mathcal{H}^{\circ}\oplus H$ where $H$ is a plane in $S^{2}_{h}V$, which in turns corresponds to a geodesic in $\mathbb{H}^{3}$ in the projective Klein model.
@@
If $S$ is a surface in $\mathbb{H}^{3}$ with principal curvature in $(-1,1)$, then the set of normal geodesics forms spacelike surface, for the pseudo-Riemannian structure on the space of geodesics described in Section A.1. The corresponding map $\mathcal{G}u:S\to\mathcal{G}$ is called the Gauss map. Nearly Fuchsian representations are representations of a closed surface group $\Gamma_{g}$ admitting an equivariant surface with principal curvature in $(-1,1)$. They are a priori a larger class of representations than almost Fuchsian representations, for which the equivariant surface with principal curvature in $(-1,1)$ is required to be minimal.
The space $\mathcal{G}$ also admits a special $\operatorname{SL}_{2}(\mathbb{C})$-invariant symplectic structure. An immersion in $\mathcal{G}$ is locally the Gauss map of an immersion with principal curvature in $(-1,1)$ if and only if it is spacelike and Lagrangian for this symplectic structure [ES22]. Therefore if the fitting immersion is not Lagrangian, it does not come as the Gauss map of a surface in $\mathbb{H}^{3}$. Hence there could be representations admitting fitting immersions that are not nearly Fuchsian.
@@
Here $V=\mathbb{C}^{2}$ and if $\rho:\Gamma_{g}\to\operatorname{SL}(2,\mathbb{C})\subset\operatorname{Sp}(4,\mathbb{R})$, the Guichard-Wienhard domain of discontinuity corresponds to the pullback in $\mathbb{RP}^{3}$ of the complement in $\mathbb{CP}^{1}$ of the limit set of $\rho$. Since $\Gamma_{g}$ is a surface group, this domain in $\mathbb{CP}^{1}$ is the union of two topological disks, and hence for each $x\in\widetilde{S_{g}}$ the geodesic corresponding to $u(x)$ has one endpoint in each of these discs.
@@
Given any $\mathcal{C}^{1}$ embedded circle $\gamma$ in $\mathbb{CP}^{1}$, and any $\epsilon>0$, there exist a quasi-Fuchsian representation $\rho:\Gamma_{g}\to\operatorname{SL}(2,\mathbb{C})$ for some genus $g$ large enough whose limit set has Haussdorf distance at most $\epsilon$ with $\gamma$.
@@
For all $t\in\mathbb{R}$, let $x_{t}\in\mathbb{CP}^{1}$ be the endpoint of the geodesic corresponding to $u(\gamma(t))$ that belongs to the interior of the Jordan curve $\Lambda$. When $t$ goes to +$\infty$+ ∞, $x_{t}$ converges to $z$ and it converges to $x$ for $t$ going to $\infty$- ∞. Moreover $x_{t}$ always belong to the circle determined by $\Phi_{t}(q)$.
@@
Appendix B Fitting immersions and spacelike immersions for $\operatorname{PSp}(4,\mathbb{R})$
In this section we explain how the data of a maximal immersion into the Pseudo-Riemannian space $\mathbb{H}^{2,2}$ with principal curvature in $(-1,1)$ induces a fitting immersion. Combining this with a result from Collier-Tholozan-Toulisse we show that our characterization of maximal representations can be improved in $\operatorname{Sp}(4,\mathbb{R})$.

### v2

@@ 5.2. Fibered domain of discontinuity
We first apply Theorem 5.4 which shows that $\rho$ is $\{n\}$-Anosov. Let $x\in\widetilde{S_{g}}$, we prove that the homotopy type of the boundary curve $\xi^{n}_{\rho}$ is the same as the homotopy type $\xi_{u(x)}]$ from Proposition 6.3. We consider the map $\zeta_{\infty}$ that associates to $q]\in\mathbb{S}u^{*}\mathcal{E}_{x}$ the limit of $\Phi_{t}([q])$ in $\partial\Gamma$. This map is homotopic in $\widetilde{S_{g}}\cup\partial\Gamma\setminus\{x\}$ to the map $\zeta_{1}$ that associates the projection $p(\Phi_{1})$. Hence by Lemma 4.3, $\zeta_{1}$ defines a diffeomorphism of degree $1$ between the circle $\mathbb{S}u(x)$ with its maximal orientation and the boundary $\partial\Gamma$ for the induced orientation.
The map $\xi^{n}_{\rho}\circ\zeta_{\infty}$ has the homotopy type $\xi_{u(x)}]$ associated to $u(x)$: it defines a boundary map as in Proposition 6.3. The fact that $\rho$ is maximal is then a consequence of the characterization of maximal representations from Theorem 2.5. Indeed $u(x)\in\operatorname{Gr}_{2}^{\max}(\mathcal{Q})$ implies that $\xi_{u(x)}]=n[\tau]$, and we already know that the degree of $\zeta_{\infty}$ is equal to the degree of $\zeta_{1}$ which is equal to $1$.
@@ 6.3. Construction of a fitting flow
A quadruple of Lagrangians $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$ is called maximal if each cyclic oriented subtriple is maximal.
@@
Let $(\ell_{1},\ell_{2},\ell_{3},\ell_{4})$ be a maximal quadruple of Lagrangians. The bilinear form $q_{\ell_{4},\ell_{3}}-q_{\ell_{1},\ell_{2}}$ is positive. In particular the zero set of these quadrics define two disjoint quadric hypersurfaces in $\mathbb{P}(\mathbb{R}^{2n})$.
@@
Let ${v}\in\mathbb{R}^{2n}$, it can be decomposed uniquely as ${v}={v}_{1}+{v}_{2}$ with ${v}_{1}\in\ell_{1}$ and ${v}_{2}\in\ell_{2}$. Moreover there exist some unique $x,y\in\ell_{4}$ such that ${v}_{1}=x+u_{1}(x)$ and ${v}_{2}=y+u_{2}(y)$. The vector v{v} decomposes therefore as ${v}=x+y+u_{1}(x)+u_{2}(y)$. One computes that :
@@
[…]
$\omega(x+u_{1}(x),\widetilde{u_{2}}(x+u_{1}(x))=\omega(x,\widetilde{u_{2}}(x+u_{1}(x))>0.$
However note that by definition $x+u_{1}(x)+\widetilde{u_{2}}(x+u_{1}(x))=x+u_{2}(x)$ and in particular $\widetilde{u_{2}}(x+u_{1}(x))=(u_{2}-u_{1})(x)$. Hence the symmetric bilinear form $\omega\left(\cdot,(u_{2}-u_{1})(\cdot)\right)$ is positive on $\ell_{4}$.
In particular for $y\neq 0$ the previous expression is strictly greater than the following $\omega\left(x,u_{1}(x)\right)+\omega\left(y,u_{1}(y)\right)+2\omega\left(y,u_{1}(x)\right)$, which is non-negative since $\omega\left(\cdot,u_{1}(\cdot)\right)$ is positive. In the case when $y=0$, this last inequality is strict for $x\neq 0$. Otherwise the previous inequality $\omega\left(y,(u_{2}-u_{1})(y)\right)>0$ is strict. Therefore for ${v}\neq 0$, $q_{\ell_{4},\ell_{3}}({v},{v})-q_{\ell_{1},\ell_{2}}({v},{v})>0$. ∎
@@
[…]
@@
[…]
@@
[…]
@@
Let $w\in T^{1}_{x}\widetilde{S}$ be such that $\langle w,v\rangle>0$. The following quadruple is positive $(\zeta_{Jw},\zeta_{-Jw},\zeta_{-J\phi(w)},\zeta_{J\phi(w)})$ due to the negative curvature of the metric we put on $\widetilde{S_{g}}$, see Figure 4. Hence the corresponding quadruple of Lagrangians is maximal. Therefore $q_{\phi(w)}^{\circ}-q_{w}^{\circ}$ is positive by Lemma 6.11. When $\langle w,v\rangle<0$, the following quadruple is positive $(\zeta_{-Jw},\zeta_{Jw},\zeta_{J\phi(w)},\zeta_{-J\phi(w)})$ and therefore $q_{\phi(w)}^{\circ}-q_{w}^{\circ}$ is negative. Hence :
@@
Given an immersion $h:M\to\mathbb{P}(S^{2}V^{>0})$ from a manifold $M$ of dimension $d$ we define its Gauss map $\mathcal{G}h:M\to\operatorname{Gr}_{d}(\mathcal{Q})$, that associates to $x\in M$ the pencil $P$ associated with the codimension $d$ projective subspace of $\text{Gr}_{d}(\mathcal{Q})$ orthogonal to $\mathrm{d}h(TM)$ at $h(x)$, for the $\operatorname{SL}(V)$-invariant Riemannian metric on $\mathbb{P}(S^{2}V)$.
@@
We first identify $\mathbb{P}(S^{2}V^{>0})$ with $\mathbb{P}(S^{2}(V^{*})^{>0})$ via the map $X]\mapsto[X^{-1}]$. Note as once again we view elements of $S^{2}V$ and $S^{2}(V^{*})=\mathcal{Q}$ respectively as maps $V^{*}\to V$ and $V\to V^{*}$. An immersion $h:M\to\mathbb{P}(S^{2}V^{>0})$ hence defines a dual immersion $h^{*}:M\to\mathbb{P}(S^{2}(V^{*})^{>0})$. Fixing a volume form on $V$ and $V^{*}$ allows us to lift this map to a map $\overline{h}^{*}$ into the space of elements in $S^{2}(V^{*})^{>0}$ whose corresponding map $V\to V^{*}$ has determinant $1$. We define $\mathcal{G}h(x)=\mathrm{d}\overline{h}^{*}(T_{x}M)\subset S^{2}V^{*}=\mathcal{Q}$.
@@
Let $\gamma:\mathbb{R}\to M$ be a geodesic for the metric induced by $h$. Recall that $\overline{h}(\gamma(t))$ for $t\in\mathbb{R}$ is an element of $S^{2}V$, and hence is defines a positive symmetric endomorphism of $V^{*}$ and as such it is diagonalizable is an orthonormal basis with positive coefficients. We write the representative of $h(\gamma(o))$ with determinant $1$ in a basis $(e_{i})_{i\in I}$ such that for some $\lambda^{i}\in\mathbb{R}$, for all $t\in\mathbb{R}$:
@@
Let us fix a symplectic form $\omega$ on $V=\mathbb{R}^{2n}$. Let $\mathbb{X}_{\operatorname{Sp}}$ be the subset of $\mathbb{P}(S^{2}V^{>0})$ consisting of tensors $q^{-1}]$ that are compatible with $\omega$, i.e. such that for some complex structure $J$ on $\mathbb{R}^{2n}$, the bilinear form $q+i\omega$ is a $J$-hermitian metric on $V$. Recall that $q:V\to V^{*}$ is a bilinear form, and $q^{-1}:V^{*}\to V$ is a tensor.
The space $\mathbb{X}_{\operatorname{Sp}}$ is a copy of the symmetric space associated to $\operatorname{Sp}(2n,\mathbb{R})$, which is a totally geodesic subspace of the symmetric space associated to the Lie group $\operatorname{SL}(2n,\mathbb{R})$ whose model is $\mathbb{P}(S^{2}V^{>0})$. However it is not a projective subset : the closure of the projective convex hull of $\mathbb{X}_{\operatorname{Sp}}$ in $\mathbb{P}(S^{2}V^{>0})$ is equal to $\mathbb{P}(S^{2}V^{\geq 0})$ since it contains all the extreme points of $\mathbb{P}(S^{2}V^{\geq 0})$, i.e. the rank one elements $S^{2}\mathbb{P}(\mathbb{R}^{2n})$.
Indeed, we can write matrices of the form $\text{Diag}(\lambda,1,\cdots,1,\lambda^{-1})\in\operatorname{Sp}(2n,\mathbb{R})$ for any $\lambda\in\mathbb{R}_{>0}$ and such that the first basis spans any fixed line $\ell\in\mathbb{RP}^{2n-1}$, one can observe that the closure of the orbit of any point of $\mathbb{X}_{\operatorname{Sp}}$ by these elements contains the rank one point corresponding to $\ell$. Hence the closure of $\mathbb{X}_{\operatorname{Sp}}$ contains all rank one points.
@@
The anihilator $(T\mathbb{X}_{\operatorname{Sp}})^{\circ}\subset S^{2}V$ of the tangent space to $\mathbb{X}_{\operatorname{Sp}}$ at this point, viewed as a linear subspace of $\mathcal{Q}$ can be identified with the space of following symmetric matrices where $A$ is symmetric and $B$ is antisymmetric:
@@
Suppose that the intersection is not transverse, i.e. that $q$ can be written in this form. Let $\ell^{+}$ and $\ell^{-}$ be Lagrangians on which $q$ is respectively positive and negative. Since the maximal compact $U(n)$ acts transitively on the space of Lagrangians one can assume that $\ell^{+}=\langle x_{1},\cdots,x_{n}\rangle$. Let $\ell_{0}^{-}=\langle y_{1},\cdots,y_{n}\rangle$ be its orthogonal for the fixed euclidean metric. The Lagrangian $\ell^{-}$ is transverse to $\ell^{+}$ so for some symmetric matrix $U$, one can $\ell^{-}$ as the image of :
@@
This bilinear form is regular if and only if all the $\lambda_{i}$ are non-zero. Else we can assume that $\lambda_{i}>0$ up to exchanging $x_{i}$ by $y_{i}$ and $y_{i}$ by $x_{i}$ in the symplectic basis. Then the bilinear form is positive on the Lagrangian $\langle x_{1},\cdots,x_{n}\rangle$ and negative on the Lagrangian $\langle y_{1},\cdots,y_{n}\rangle$. ∎
@@ Appendix A Immersions in the space of geodesics in $\mathbb{H}^{3}$
In this section we set $V=\mathbb{C}^{2}$, and restrict ourselves to pencils of Hermitian quadrics. Note that we can see this way $\operatorname{SL}(2,\mathbb{C})$ as a subgroup of $\operatorname{Sp}(4,\mathbb{R})\simeq\operatorname{Sp}(V,\omega)$ for $\omega$ the real part of a volume form on $\mathbb{C}^{2}$. The results of this Section serve as an illustration of the previously introduced notions of fitting pairs and fitting directions.
@@
Let $S^{2}V=S^{2}_{h}V\oplus S^{2}_{a}V$ be the eigenspace decomposition for the operator $J\otimes J$ where $J$ is the complex conjugation. Here $S^{2}_{h}V$ is the $4$-dimensional eigenspace associated to $1$ and $S^{2}_{a}V=\mathcal{H}^{\circ}$, the $6$-dimensional space associated to 1-1. The intersection of $\mathbb{P}(S^{2}_{h}V)$ with the space of positive tensors is the projective Klein model for $\mathbb{H}^{3}$.
The annihilator of a pencil $P\in\operatorname{Gr}_{2}^{{\operatorname{mix}}}(\mathcal{H})$, i.e. a plane $P\subset\mathcal{H}$ such that no non-zero vectors are semi-positive, is equal to $P^{\circ}=\mathcal{H}^{\circ}\oplus H$ where $H$ is a plane in $S^{2}_{h}V$, which in turns corresponds to a geodesic in $\mathbb{H}^{3}$ in the projective Klein model.
@@
Remark A.1.
The space $\mathcal{G}$ is a $\operatorname{SL}(2,\mathbb{C})$-homogeneous spac. Note that however the larger spaces $P\in\operatorname{Gr}^{{\operatorname{mix}}}_{2}(\mathcal{Q})$ or $P\in\operatorname{Gr}^{{\omega}}_{2}(\mathcal{Q})$ from the rest of the paper are not $\operatorname{Sp}(2n,\mathbb{R})$ or $\operatorname{SL}(2n,\mathbb{R})$ homogeneous spaces.
@@
If $S$ is a surface in $\mathbb{H}^{3}$ with principal curvature in (1,1)(-1,1), then the set of normal geodesics forms spacelike surface, for the pseudo-Riemannian structure on the space of geodesics described in Section A.1. The corresponding map $\mathcal{G}u:S\to\mathcal{G}$ is called the Gauss map. Nearly Fuchsian representations are representations of a closed surface group $\Gamma_{g}$ admitting an equivariant surface with principal curvature in (1,1)(-1,1). They are a priori a larger class of representations than almost Fuchsian representations, for which the equivariant surface with principal curvature in (1,1)(-1,1) is required to be minimal.
The space $\mathcal{G}$ also admits a special $\operatorname{SL}_{2}(\mathbb{C})$-invariant symplectic structure. An immersion in $\mathcal{G}$ is locally the Gauss map of an immersion with principal curvature in (1,1)(-1,1) if and only if it is spacelike and Lagrangian for this symplectic structure [ES22]. Therefore if the fitting immersion is not Lagrangian, it does not come as the Gauss map of a surface in $\mathbb{H}^{3}$. Hence there could be representations admitting fitting immersions that are not nearly Fuchsian.
@@
A particular corollary is the following result on Nearly Fuchsian representations, introduced in Remark A.4.
Corollary A.6.
There exist quasi-Fuchsian representations that are not nearly Fuchsian.
Almost Fuchsian representations were already known to not contain all quasi-Fuchsian representations see for isntance [HW15], but a recent result of Nguyen, Schlenker and Seppi [NSS25] proves that nearly Fuchsian representations are not all almost Fuchsian, making this corollary new.
Here $V=\mathbb{C}^{2}$ and if $\rho:\Gamma_{g}\to\operatorname{SL}(2,\mathbb{C})\subset\operatorname{Sp}(4,\mathbb{R})$, the Guichard-Wienhard domain of discontinuity $\Omega_{\rho}$, mentionned in Proposition 5.5 corresponds to the pullback in $\mathbb{RP}^{3}=\mathbb{P}_{\mathbb{R}}(\mathbb{C}^{2})$, of a domain of discontinuity $\Omega_{\rho}^{0}\subset\mathbb{CP}^{1}=\mathbb{P}_{\mathbb{C}}(\mathbb{C}^{2})$. This domain is the complement in $\mathbb{CP}^{1}$ of the limit set of $\rho$.
Since $\Gamma_{g}$ is a surface group, this domain in $\mathbb{CP}^{1}$ is the union of two topological disks, and hence for each $x\in\widetilde{S_{g}}$ the geodesic corresponding to $u(x)$ has one endpoint in each of these discs. This fact will be used in the proof of Theorem A.5.
@@
Given any $\mathcal{C}^{1}$ embedded circle $\gamma$ in $\mathbb{CP}^{1}$, and any $\epsilon>0$, there exist a quasi-Fuchsian representation $\rho:\Gamma_{g}\to\operatorname{SL}(2,\mathbb{C})$ for some genus $g$ large enough whose limit set has Hausdorff distance at most $\epsilon$ with $\gamma$.
@@
For all $t\in\mathbb{R}$, let $x_{t}\in\mathbb{CP}^{1}$ be the endpoint of the geodesic corresponding to $u(\gamma(t))$ that belongs to the interior of the Jordan curve $\Lambda$. When $t$ goes to $\infty$, $x_{t}$ converges to $z$ and it converges to $x$ for $t$ going to -$\infty$. Moreover $x_{t}$ always belong to the circle determined by $\Phi_{t}(q)$.
@@
Appendix B Fitting immersions and spacelike immersions for $\operatorname{Sp}(4,\mathbb{R})$
In this section we explain how the data of a maximal immersion into the Pseudo-Riemannian space $\mathbb{H}^{2,2}$ with principal curvature in (1,1)(-1,1) induces a fitting immersion. Combining this with a result from Collier-Tholozan-Toulisse we show that our characterization of maximal representations can be improved in $\operatorname{Sp}(4,\mathbb{R})$.